A normally distributed population has a mean of 60 and a standard deviation of 6.
Compute the probability of a value between 51 and 66. ________
Compute the probability of a value less than 51. ________
Compute the probability of a value greater than 72. ________

Answer by stanbon(75887)

A normally distributed population has a mean of 60 and a stand deviation of 6.
-------------
You need to find the z-value on each of these:
For example:
Compute the probability of a value between 51 and 66. ________
z(66) = (66-60)/6 = 1
z(51) = (51-60)/6 = -1.5
P(51 < x < 66) = P(-1.5 < z < 1) = 0.77454..
-----------------------------------------------

Compute the probability of a value less than 51. ________
P(z < 51) = P(z< -1.5) = 0.06681..
-----------------------------------------------

Compute the probability of a value greater than 72. ________
z(72) = (72 - 60)/6 = 12/6 = 2
P(x > 72) = P(z > 2) = 0.02275...
-------------------
Cheers,
Stan H.